You and 11 coworkers just won $14 million ($1,166,666.67 each) from the state lottery. Assuming you each receive your share over 17 years and that the state lottery earns a 8 percent return on its funds, what is the present value of your prize before taxes if you request the up-front cash option? Click on the table icon to view the annuity table The present value of your prize before taxes if you request the up-front cash option is S Round your answer to the nearest dollar.)
Answer #1


solution: PV=FV/(1+i)^n PV=1166666.67/(1+0.08)^17 PV= FV Future value PV Present value i rate of return per period n Number o